ms. burke invested $27,000 in two account one yeilding 7% interest and the other 8%. If she received a total of $2,060 in interest at the end of the year, how much did she invest in each account?
Answer by mananth(16949)

Investment I 7.00% per annum x
Investment II 8.00% per annum y

x+y=27000 ------------------------1
7.00%x+ 8.00%y= $2,060.00
Multiply by 100
7x+8y=$206,000.00 --------2

Multiply (1) by -7
we get

-7x-7y= -189000.00

Add this to (2)

y = $17,000.00 investment at 8.00%
Balance $10,000.00 investment at 7.00%
CHECK
$10,000.00 $700.00
$17,000.00 $1,360.00
Total $2,060.00